OpenStudy (anonymous):

I know the answer, just need to learn how to solve it Consider the following. g(x)=x^3-64/x^224 Find the zeros (if any) of the rational function. (Enter NONE in any unused answer blanks.) Use a graphing utility to verify your answer.

OpenStudy (anonymous):

answer are 4 and none.

OpenStudy (amistre64):

when the top = 0 we got a 0 function

OpenStudy (amistre64):

so long as the bottom doesnt get a bad result

OpenStudy (anonymous):

so you replace x by 0?

OpenStudy (amistre64):

x^3 -64 = 0

OpenStudy (amistre64):

factor the diff of cubes

OpenStudy (anonymous):

(x-4)^3

OpenStudy (amistre64):

(x-4)(x^2+4x+16) = 0

OpenStudy (anonymous):

is what i wrote correct/?

OpenStudy (amistre64):

no; you need to factor the 'difference of cubes' not make up a cubic expression from the equation :)

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

ok, so le't get going from your expression

OpenStudy (amistre64):

the right part is irreducible so it doesnt factor or have zeros the left part is: x-4 = 0, x=4

OpenStudy (anonymous):

wait, how you get x-4=0 and x=4?

OpenStudy (amistre64):

\[(x^3-64)\] factor the difference of cubes: \[(x-4)(x^2+4x+16)\] solve for zeros since 0(B) = 0 or A(0) = 0

OpenStudy (anonymous):

oh i totally forogt about the

OpenStudy (amistre64):

\[x-4=0 \implies x=4\]

OpenStudy (amistre64):

\[x^2+4x+16 =0 \implies never\]

OpenStudy (anonymous):

how you get zero?

OpenStudy (amistre64):

Do you mean; How do I determine when each factor equals zero?

OpenStudy (anonymous):

yeha

OpenStudy (amistre64):

i set each factor equal to zero and solve for x

OpenStudy (amistre64):

x-4 =0 when ???? x-4 = 0 ; +4 each side x-4+4 = 0+4 x-0 = 4 x=4

OpenStudy (anonymous):

i understand how you get 4, but i dont get how you get 0

OpenStudy (amistre64):

soosfgosgoasogosrgoi.......

OpenStudy (amistre64):

ok.... tell me, does: \[0*anything=0?\]

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

i got another problem simillar to this one

OpenStudy (amistre64):

:) what is it

OpenStudy (anonymous):

g(x)=x^3-1/x^2+6

OpenStudy (anonymous):

same question

OpenStudy (amistre64):

x^3-1 ------ = 0 when the top =0 right? x^2+6

OpenStudy (anonymous):

yeah

OpenStudy (amistre64):

factor the difference of cubes on the top to get: (x-1) (x^2+x+1)

OpenStudy (anonymous):

you get 1, and 0

OpenStudy (amistre64):

we know that: 0 * (x^2 +x+1) = 0 and that: (x-1) *0 = 0

OpenStudy (anonymous):

thx man, for the help

OpenStudy (amistre64):

x-1 = 0; when x=1 x^2+x+1; has no zero value

OpenStudy (amistre64):

youre welcome :)

OpenStudy (anonymous):

\[x = 2^{(6/227)} \]

OpenStudy (anonymous):

A plot supporting the conclusion that x=2^(6/227) is attached.